# GNN: Graph Neural Network and Large Language Model Based for Data Discovery

### v1

@@ Abstract
Our algorithm GNN: Graph Neural Network and Large Language Model Based for Data Discovery inherits the benefits of [Hoang(2024b)] (PLOD: Predictive Learning Optimal Data Discovery), [Hoang(2024a)] (BOD: Blindly Optimal Data Discovery) in terms of overcoming the challenges of having to predefine utility function and the human input for attribute ranking, which helps prevent the time-consuming loop process. In addition to these previous works, our algorithm GNN leverages the advantages of graph neural networks and large language models to understand text type values that cannot be understood by PLOD and MOD, thus making the task of predicting outcomes more reliable. GNN could be seen as an extension of PLOD in terms of understanding the text type value and the user’s preferences based on not only numerical values but also text values, making the promise of data science and analytics purposes.
@@ IV-A Runtime Comparison
We show in the figure LABEL:Runtime_comparison_with_changes_in_the_number_of_tuples that all algorithms’ runtime increases linearly with the number of tuples. Among the algorithms, the GNN exhibited a slightly lower runtime than others. The BOD algorithm demonstrated a relatively higher runtime due to its complex decision-making process. In contrast, the Multi-objective and Skyline algorithms performed comparably, showing similar runtime patterns, while the Top-K algorithm had a runtime close to PLOD.

### v2

# GNN: Graph Neural Network and Large Language Model for Data Discovery

@@ Abstract
Our algorithm GNN: Graph Neural Network and Large Language Model for Data Discovery inherit the benefits of [Hoang(2024b)] (PLOD: Predictive Learning Optimal Data Discovery), [Hoang(2024a)] (BOD: Blindly Optimal Data Discovery) in terms of overcoming the challenges of having to predefine utility function and the human input for attribute ranking, which helps prevent the time-consuming loop process. In addition to these previous works, our algorithm GNN leverages the advantages of graph neural networks and large language models to understand text type values that cannot be understood by PLOD and MOD, thus making the task of predicting outcomes more reliable. GNN could be seen as an extension of PLOD in terms of understanding the text type value and the user’s preferences based on not only numerical values but also text values, making the promise of data science and analytics purposes.
@@ IV-A Runtime Comparison
We show in the figure 2 that all algorithms’ runtime increases linearly with the number of tuples. Among the algorithms, the GNN exhibited a slightly lower runtime than others. The BOD algorithm demonstrated a relatively higher runtime due to its complex decision-making process. In contrast, the Multi-objective and Skyline algorithms performed comparably, showing similar runtime patterns, while the Top-K algorithm had a runtime close to PLOD.